OpenStudy (anonymous):

A swimming pool has to be drained for maintenance. The pool is shaped like a cylinder with a diameter of 15m and a depth of 1.7m . If the water is pumped out of the pool at the rate of 17m^3 per hour, how many hours does it take to empty the pool? Use the value 3.14 for PIE , and round your answer to the nearest hour.

OpenStudy (anonymous):

frist find Volume of the poor V= A*H V =15^2*pi*1.7 V = 1,201,65919 \[\frac{1h}{h?}=\frac{17}{1,201,65919}\]Cros and time\[\frac{1,201,65919}{17}=70.6858\] 70.7 hours

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

no thank u 4 another great explanation ;-)

OpenStudy (anonymous):

u r w come